OpenStudy (anonymous):

Simplify the product. 5b(6b - 10) A. 11b - 10 B. 5b2 + 6b - 10 C. -b - 10 D. 30b2 - 50b

pooja195 (pooja195):

Any ideas on where to start?

OpenStudy (anonymous):

no idea at all

pooja195 (pooja195):

Start off by distributing 

OpenStudy (anonymous):

11b(-10) ?

pooja195 (pooja195):

nope we multiply 5 x 6 =?

OpenStudy (anonymous):

30b(-10) ?

pooja195 (pooja195):

b*b=b^2 so 30b^2 5b * -10 =?

OpenStudy (anonymous):

-50b ?

pooja195 (pooja195):

perfect! :) so our answer is...?

OpenStudy (anonymous):

i don't know im sorry

pooja195 (pooja195):

we have 30b^2 and -50b put them together :)

TheSmartOne (thesmartone):

Pooja didn't quite tell you what the distributive property is: a(b-c) = ab - ac

OpenStudy (anonymous):

ohhhh 30b^2 * -50b

pooja195 (pooja195):

Nope we don't add the multiplcation sign remove it now what do we get?

OpenStudy (anonymous):

30b^2 -50b ?

pooja195 (pooja195):

perfect :)

OpenStudy (anonymous):

thank you!!!!!

pooja195 (pooja195):

yw ^_^